Calculate the concentrations of H +, OH-, pH, pOH of aqueous sulfuric acid solution with a concentration of 0.0049 mol / dm3 assuming complete dissociation!


1
Expert's answer
2021-05-21T02:33:41-0400

Since sulfuric acid is a strong acid, H2SO4 dissociates completely. The molar concentration of protons in the solution equals:

[H+] = 0.0049 M × 2 = 0.0098 M

pH = -log[H+] = -log[0.0098] = 2

p[OH] = 14 - pH = 14 - 2 = 12

[OH-] = 10-pOH = 10-12 M


Answer: 0.0098 M; 10-12 M; 2; 12